OpenStudy (anonymous):

The length of a field is three times its width. If the perimeter is 64 m, what is the width of the field? A. 8 m B. 16 m C. 21 m D. 32 m

OpenStudy (anonymous):

who ever helps gets a medal

OpenStudy (anonymous):

if you help i will give you a medal

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

I will help just give me a couple sec.

OpenStudy (anonymous):

thanxs

OpenStudy (anonymous):

A.8

OpenStudy (anonymous):

thanks here is your medal

OpenStudy (anonymous):

The answer is A Write: L =3w 2L+2w=64 2(3w)+2w =64 6w+2w=64 8w=64 w=64/8 w=8

OpenStudy (anonymous):

Go and Google perimeter of a rectangle and then go through each choice ex 8 for width then do 8 *3 = 24 for lenth and it equals 64